3(x)-2(y)=12
can only use the number 5, 6, 7, 8, or 9

Answer by Edwin McCravy(20054)


3(x)-2(y)=12
can only use the number 5, 6, 7, 8, or 9

---------------------------------------

With a little trial and error we get

x = 8 and y = 6

because

3(8)-2(6) = 24-12 = 12

Edwin